The following data resulted from 5 experimental runa made on four indrpendent variables and a single response y


1
Expert's answer
2021-09-21T09:55:52-0400

Since given data is incomplete, we assume it as follows:

The following data resulted from 5 experimental runs made on four independent variables and a single response y.



Here n=5. Write 5 equations with 4 explanatory variables in matrix notation.

Solution:

Here n=5, k=4.

Assuming that the model is

 "y=\beta_{0}+\beta_{1} X_{1}+\beta_{2} X_{2}+\ldots+\beta_{4} X_{4}+\varepsilon"  

the n-tuples of observations are also assumed to follow the same model. Thus they satisfy

"\begin{aligned}
&y_{1}=\beta_{0}+\beta_{1} x_{11}+\beta_{2} x_{12}+\ldots+\beta_{4} x_{1 4}+\varepsilon_{1} \\

&\vdots \\
&y_{4}=\beta_{0}+\beta_{1} x_{5 1}+\beta_{2} x_{5 2}+\ldots+\beta_{4} x_{54}+\varepsilon_{5}
\end{aligned}"

These 5 equations can be written in matrix notation as

"\left(\begin{array}{l}y_{1} \\ y_{2} \\ \vdots \\ y_{5}\end{array}\right)=\left(\begin{array}{ccccc}1 & x_{11} & x_{12} & \cdots & x_{1 4} \\ 1 & x_{21} & x_{22} & \cdots & x_{2 4} \\ \vdots & \vdots & \vdots & \vdots \\ 1 & x_{5 1} & x_{5 2} & \cdots & x_{54}\end{array}\right)\left(\begin{array}{l}\beta_{0} \\ \beta_{1} \\ \vdots \\ \beta_{4}\end{array}\right)+\left(\begin{array}{l}\varepsilon_{1} \\ \varepsilon_{2} \\ \vdots \\ \varepsilon_{5}\end{array}\right)"